Flying against the wind, an airplane travels 4200 km in 5 hours. Flying with the wind, the same plane travels 7840 km in 7 hours. What is the rate of the plane in still air and what is the rate of the wind. Show work if you can so I can learn how to work it out. Please and thanks

Answer by TimothyLamb(4379)

x = airplane speed in still air
y = wind speed
s = d/t
---
against wind:
x - y = 4200/5
---
with the wind:
x + y = 7840/7
---
put the system of linear equations into standard form
---
x - y = 4200/5
x + y = 7840/7
---
x - y = 840
x + y = 1120
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = airplane speed in still air = 980 kph
y = wind speed = 140 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php